OpenStudy (anonymous):

PLEEEEAAAASE HELP?? D: Solve. 4(a + 2) = 14 – 2(3 – 2a) A. –1 B. –2 C. all real numbers D. no solution

OpenStudy (arthur_ser):

4(a+2)=14-2(3-2a) 4a+8=14-6+4a 14=14 Thus the correct answer is C

OpenStudy (lovelyanna):

I agree, C is correct

OpenStudy (anonymous):

thank you both.